Circular No. 7353

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000F IN IC 302 M. Papenkova, and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126) report the discovery of an apparent supernova on unfiltered images taken on Jan. 29.2 and 30.2 (both with mag ∼ 17.5) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2000F is located at α = 3^{h}12^{m}52^{s}.71, δ = +4^{o}42′34′′.4 (equinox 2000.0), which is 21′′.2 east and 8′′.8 north of the nucleus of IC 302. An unfiltered KAIT image taken on Jan. 10.2 already showed a hint of the new object, while an image taken on 1999 Dec. 30.3 showed nothing at the position of SN 2000F (limiting mag 19.0).

SUPERNOVA 2000E IN NGC 6951 V. F. Polcaro, CNR-IAS, Rome; and A. de Blasi, R. Gualandi, and

S. Galleti, Bologna Astronomical Observatory (BAO), report: “Low- and intermediate-resolution spectra of SN 2000E (cf. IAUC 7351) were taken with the 1.52-m telescope (+ BFOSC) at Loiano on Jan. 28 and 30. These spectra clearly show a narrow absorption feature at 627.8 ± 0.3 nm. This feature is compatible, inside the instrumental uncertainties, with an Hα absorption at the host galaxy redshift. If this hypothesis is true and if SN2000E will behave as did type-Ia SN 1990M, the feature will disappear in a few days. Further observations are urgently needed.”

Further to the announcement of SN 2000E on IAUC 7351, E. Brocato adds some additional information. On line 5, for α = 20^{h}37^{m}13^{s}.8 read α = 20^{h}37^{m}13^{s}.77. On Jan. 25.737 UT, J = 14.5 ± 0.1. The limiting magnitudes for the images that did not show SN 2000E are as follows: Digital Sky Survey, ∼ 21; 1999 Dec. 18.78, V = 18.0; 2000 Jan. 6.74, J = 19.16, H = 19.38.

IY URSAE MAJORIS N. N. Samus, Sternberg Astronomical Institute, Moscow, on behalf of

IAU Commission 27, informs us that the new cataclysmic variable reported by Uemura et al. on IAUC 7349 has been given the designation IY UMa.

SUPERNOVA 2000B IN NGC 2320 Unfiltered CCD magnitudes: Jan. 21.194 UT, 16.70 (M. Suárez, Las

Palmas, Canary Islands); 25.719, 15.8 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan).

2000 January 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green